OpenStudy (jetman24):

Will give medal to whoever solves this: True or False: 1/x5 = x-5

OpenStudy (curry):

this is a wrong question. It's a quadratic equation with the determinant > 0. It should really be what are the solutions. But if it really does ask T/F, then it's obvi T, as you will have 2 different solutions.

OpenStudy (curry):

5x^2 - 25x - 1 = 0 would be the quadratic equation.

OpenStudy (mrnood):

this is not a quadratic - as it has a power of 5 in it @jetman24 - the question is not complete - it does not have a T/F answer as you have written it

OpenStudy (mrnood):

@Curry Soz - I assume x5 MEANS x^5

OpenStudy (curry):

I think he meant 5x, but wrote it the other way around.

OpenStudy (curry):

thought*

OpenStudy (mrnood):

Either way - it does not have a T/F answer

OpenStudy (curry):

Yep, my point exactly.

OpenStudy (jetman24):

@Curry @MrNood I mean 1/x^5 = x-5

OpenStudy (jetman24):

that is the true equation

OpenStudy (jetman24):

Can you please solve it?

OpenStudy (curry):

2 real and 4 complex solutions. Kinda really difficult to solve by hand.

OpenStudy (mrnood):

Do you have a value of x to solve it for? That is the only way you would get a True/False answer